# Exceptional topology in Non-Hermitian Twisted Bilayer Graphene

### v1

@@ Abstract
Twisted bilayer graphene has extraordinary electronic properties at the magic angle along with an isolated flat band at magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of Bistritzer-MacDonald(BM) model, we find Dirac cones centered at corner of the moiré Brillouin zone deformed in the presence of non-Hermiticity. This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both and corners of the Brillouion zone. The coincident of exceptional rings at point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime. More interestingly, we find that the topological charge in the moiré Brillouin zone is conserved during the expansion and fusion of the exceptional ring, which is absent in two-dimensional systems constraining by Nielsen-Ninomiya theorem. These findings can be demonstrated in realistic cold atom and metamaterial systems and will stimulate further study on non-Hermitian phenomena in twistronic.
@@
Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its wave-function topology. In single-layer graphene, the presence of unusual linear band dispersion termed Dirac band dispersion at and points of the Brillouin zone is derived from its honeycomb lattice geometry. The Lorentz invariant Dirac cone at and points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4]. When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phrases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29].
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [30, 31, 32]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or Weyl point, into a ring of exceptional points [33, 34]. At an exceptional point, not only the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including non-Hermitian skin effect [35] and non-Hermitian topological classifications [36, 37].
Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [38, 39].
@@
In particular, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in 2D periodic systems following the Nielsen-Ninomiya theorem [40, 41]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges. Actually, when the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [34]. It is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
@@
In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers. As shown in Fig. 1(a), balanced gain (in bright color) and loss (in light color) are put on the bottom layers (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the and corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig. 1(c). We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings, this is a large difference from non-Hermitian single-layer graphene.
@@ II The non-Hermitian bilayer model
We consider balanced gain and loss on both layers of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ is:
@@
Here, is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity. And corresponds to the rotated Pauli matrices . is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
@@ III The appearance and development of exceptional rings
Let us investigate the band structure for twist angle . In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent (), the dispersion in the moiré Brillouin zone is equivalent to the folded dispersion in the Brillouin zone of graphene. The Dirac point of the point in layer 1 is mapped to , while that of the point of layer 2 is mapped to .
In the presence of non-Hermitian perturbation , the energy dispersion becomes complex. In Fig. 2(a), we can see that non-Hermitian perturbation deforms the Dirac cones at corners. More specifically, the degeneracy point at the points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with onsite gain and loss. The latter has deformed Dirac cones in both and corners in the Brillouin zone. This difference is due to the fact that and in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian can only deform the Dirac cone on the bottom layer.
@@
At a critical non-Hermitian strength, the exceptional point approach point, where the whole lowest band becomes flat, as shown in Fig. 3. Different metric including the vanishing of Dirac velocity, maximum gap to neighboring bands, and minimum bandwidth gives the same results. The value of this critical non-Hermiticity strength depends on the twist angle .
@@ IV Exceptional cone
To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. LABEL:eq:H at the first honeycomb shell. The Hamiltonian is written as
@@
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in AB region . To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases. When , increases with increasing interlayer coupling in AB region . Thus, the ratio becomes larger as increases. When, increases with increasing interlayer coupling in AA region . The enon-Hermitian effect is more complicated.
@@ V Berry Curvature and Chern number
Upon summing the Berry curvature on the mmBZ, we can obtain the total Berry charge or Chern number
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. We find that the total Berry charge in mBZ is 1 throughout the merging of exceptional rings with increasing (e.g., for ). At much larger , the Berry charge can be a quantized value larger than 1 due to approaching band crossings with the remote bands. This supports our expectation that the topological charge in the system is preserved in the exceptional contour. And most importantly, this characterizes the topological features of non-Hermitian TBG. This is opposite to the merging of exceptional contours with opposite topological charge, in which the topological charge is dissipated and gives a single, uncharged exceptional contour.
@@ VI Discussion and conclusion
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings centered at corner in the moiré Brillouin zone in the () valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This flat band is characterized with a robust Chern number in moiré Brillouin zone. This is different from single-layer graphene in which Dirac/Weyl exceptional rings carrying opposite topological charges are coupled and annihilated.
@@
Note added-. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed.
@@ Acknowledgements.
- Neto et al. [2009] A. C. Neto, F. Guinea, N. M. Peres, K. S. Novoselov, and A. K. Geim, The electronic properties of graphene, Reviews of modern physics 81, 109 (2009).
@@
- McCann [2006] E. McCann, Asymmetry gap in the electronic band structure of bilayer graphene, Phys. Rev. B 74, 161403 (2006).
@@
- Bistritzer and MacDonald [2011b] R. Bistritzer and A. MacDonald, Moiré butterflies in twisted bilayer graphene, Physical Review B 84, 035440 (2011b).
@@
- Luican et al. [2011] A. Luican, G. Li, A. Reina, J. Kong, R. Nair, K. S. Novoselov, A. K. Geim, and E. Andrei, Single-layer behavior and its breakdown in twisted graphene layers, Physical review letters 106, 126802 (2011).
@@
- Shao et al. [2024] K. Shao, H. Geng, E. Liu, J. L. Lado, W. Chen, and D. Xing, Non-hermitian moiré valley filter, Physical Review Letters 132, 156301 (2024).
@@
- Tarnopolsky et al. [2019a] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Physical review letters 122, 106405 (2019a).
- Nam and Koshino [2017] N. N. T. Nam and M. Koshino, Lattice relaxation and energy band modulation in twisted bilayer graphene, Phys. Rev. B 96, 075311 (2017).
- Tarnopolsky et al. [2019b] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Phys. Rev. Lett. 122, 106405 (2019b).

### v2

@@ Abstract
Twisted bilayer graphene has extraordinary electronic properties at the magic angle along with an isolated flat band at magic angle. However, the non-Hermitian phenomena in twisted bilayer graphene remain unexplored. In this work, we study a non-Hermitian TBG formed by one-layer graphene twisted relative to another layer with gain and loss. Using a non-Hermitian generalization of Bistritzer-MacDonald model, we find Dirac cones centered at only () corner of the moiré Brillouin zone at () valley deformed in the presence of non-Hermiticity. This is different from single layer graphene with gain and loss, where rings of exceptional points appear in both and corners of the Brillouin zone. The coincident of exceptional rings at point characterizes an “exceptional magic angle”, at which the system hosts flat bands with zero energy and finite lifetime. More interestingly, we find that the topological charge in the moiré Brillouin zone is conserved during the expansion and fusion of the exceptional ring, which is absent in two-dimensional systems constraining by Nielsen-Ninomiya theorem. These findings can be demonstrated in realistic cold atom and metamaterial systems and will stimulate further study on non-Hermitian phenomena in twistronic.
@@
Graphene owes extraordinary electronic and optical properties in two-dimensional (2D) systems [1, 2]. These properties relate to its lattice structure, which is directly interconnected with its topological property. In single-layer graphene, the presence of unusual linear band dispersion termed “Dirac cone” at and points of the Brillouin zone is derived from its honeycomb lattice geometry. The two Dirac points are related by time-reversal symmetry and thus have opposite Berry phases, which has been confirmed in experiments [3, 4].
@@
When two graphene layers are shifted with each other in rational competing periodicities upon shearing or twisting, they form a moiré pattern [5, 6]. One prominent example is a graphene bilayer with a relatively small angle rotation between the layers [7]. The recent studies on twisted bilayer graphene (TBG) focus on the extraordinary electronic properties at the magic angle along with an isolated flat band [8, 9, 10, 11, 12]. It is related to a variety of exotic phases including quantum anomalous Hall [13, 14, 15, 16, 17], fractional Chern insulators [18, 19, 20], ferromagnetic states [14, 21, 22], nematicity [23], unconventional superconductivity [24, 25, 26, 27, 28, 29]. On the other hand, twisted bilayer graphene has extraordinary topological properties. Since the two Dirac points in moiré Brillouin zone (mBZ) in one valley emanate from different layers, their Berry phases can be identical due to the symmetry interlayer hopping term, which is generally not allowed in two-dimensional periodic systems following the Nielsen-Ninomiya theorem [30, 31]. This is in contrast to single-layer graphene, in which the Berry phases of Dirac cones in different valleys has different topological charges.
The topological properties of two-dimensional systems can be generalized to non-Hermitian systems by introducing on-site gain and loss [32, 33, 34]. The presence of non-Hermiticity can transform Hermitian degenerate points, such as a Dirac-like point or Weyl point, into a ring of exceptional points [35, 36]. At an exceptional point, not only the real and imaginary parts of the eigenvalues degenerate, but also the eigenvectors coalesce. Many phenomena are unique to non-Hermitian systems, including non-Hermitian skin effect [37] and non-Hermitian topological classifications [38, 39]. Despite topological robustness has been shown in one-dimensional moiré lattice under stain, the non-Hermitian effect of twisted moiré system with magic angle has been overlooked [40, 41].
Usually, non-Hermiticity is detrimental to the topology in single-layer graphene. When the exceptional contours with opposite topological charges merge, the topological charge can be dissipated [36]. In single-layer graphene, the annihilation of exceptional points from different valleys has been experimentally observed in a two-dimensional photonic system [42]. Since the moire BZ valley of TBG has different topological properties from single-layer graphene, it is natural to ask how the exceptional geometry appears and develops in a non-Hermitian TBG system.
@@
In this paper, we uncover a new type of twisted band engineering in TBG, with non-Hermitian perturbation on one of the layers. As shown in Fig. 1(a), balanced gain and loss are put on the bottom layers (in red and blue, respectively). The superposing of two graphene layers with a twist angle between them creates moiré patterns. In the mBZ shown in Fig. 1(b), the low-energy band structure consists of Dirac cones from the rotated layers located at the and corners. When the non-Hermitian perturbation is turned on the bottom layer, Dirac points at () corner in the mBZ at () valley morph into exceptional rings and the corresponding Dirac cones become exceptional cones, as shown in Fig. 1(c) and (d) respectively. We find that this non-Hermitian system hosts special angles for which multiple exceptional cones coincide at points in an mBZ. We call this phenomenon “exceptional magic”. And importantly, our results demonstrated that the total Berry charge is conserved in the mBZ even after the merging of exceptional rings, this is a large difference from non-Hermitian single-layer graphene.
@@ II The non-Hermitian bilayer model
We consider balanced gain and loss on the bottom layer of the bilayer system. We can write down a non-Hermitian generalization of the single-valley model for twisted bilayer graphene. The low-energy Hamiltonian for the mBZ at or valley is:
@@
Here, is a balanced gain and loss imposing on the bottom layer, which induces non-Hermiticity. And corresponds to the rotated Pauli matrices with for and valley respectively. is the sublattice-dependent moiré potential that couples the two layers. The interlayer coupling is in the form
@@ III The appearance and development of exceptional rings
Let us investigate the band structure for twist angle . In the absence of non-Hermiticity, this is not a magic angle with a flat band. If the interlayer coupling effect is absent (), the dispersion in the mBZ is equivalent to the folded dispersion in the Brillouin zone of graphene. Since the physics in valley and valley are related, we only discuss the valley in the following. The Dirac point at the corner in the top layer is mapped to point in the mBZ, while that at the corner of the bottom layer is mapped to point in the mBZ.
In the presence of non-Hermitian perturbation , the energy dispersion becomes complex. In Fig. 2(a), we can see that non-Hermitian perturbation deforms the Dirac cones at corners. More specifically, the degeneracy point at the points in the real part of the spectrum morphs into a ring of degeneracy points. This is a stark difference from single-layer graphene with onsite gain and loss. The latter has deformed Dirac cones in both and corners in the Brillouin zone. This difference is due to the fact that the Dirac cones at and point in non-Hermitian TBG are folded from different graphene layers. The non-Hermitian can only deform the Dirac cone on the bottom layer.
@@
At a critical non-Hermitian strength , the exceptional point approach point, where the real part whole lowest band becomes flat, as shown in Fig. 3. This characterizes an “exceptional magic” angle. According to the geometry relationship, is determined by the lattice size of mBZ and Fermi velocity .
@@ IV Exceptional cone
To understand the nonhermitian physics, we use the simplified tripodmode model [8], which truncates Eq. 1 at the first honeycomb shell. The Hamiltonian is written as
@@
We can see that aside from a renormalized velocity and non-Hermitian strength, the form of Hamiltonian is identical to the continuum model Hamiltonian of non-Hermitian single-layer graphene. The form of the renormalized velocity is the same as that of the Hermitian TBG. It decreases with increasing interlayer coupling in AB region . To see the effect of interlayer coupling on non-Hermitian strength, we should discuss two cases for and . When , increases with increasing interlayer coupling in AB region . Thus, the ratio becomes larger as increases. When , increases with increasing interlayer coupling in AA region and the non-Hermitian effect is more complicated. We can expect that the exceptional magic angle is not only determined by but also .
@@ V Berry Curvature and Chern number
Upon summing the Berry curvature on the mBZ, we can obtain the total Berry charge or Chern number
The non-Hermitian perturbation transforms a Dirac point into an exceptional ring. In the mBZ, there are three-thirds of exceptional rings from Dirac points in different moiré corners. For small , we can find that the total Berry charge in one mBZ is 1. This can be easy to understand from two facts. One is that each of the three Dirac points corresponds to Berry curvature monopoles of charge 1 in the absence of non-Hermiticity. The other is that the topological charge is preserved on the exceptional contour that forms from the original Dirac point. Thus, we can expect that each of the three exceptional contours contributes one-thirds of the Berry charge to one mBZ.
Most interestingly, even when exceeds , the total Berry charge remains one. This is opposite to the merging of exceptional contours with opposite topological charge [36], in which the topological charge is dissipated and gives a single, uncharged exceptional contour. This indicates a breakdown of Nielsen-Ninomiya theorem and distinguishes the valley topological features of non-Hermitian TBG from a single-layer graphene.
At much larger (e.g., for ), the Berry charge can be a quantized value larger than 1 due to the band crossings with the remote bands.
@@ VI Discussion and conclusion
In this paper, we have studied a non-Hermitian generalization of twisted bilayer graphene system with balanced gain and loss on one of the layers. We found that exceptional rings centered at () corner in the moiré Brillouin zone in the () valley. This is due to the folding of the Brillouin zone, and it is absent in the single-layer graphene. The exceptional rings coincide and form bands whose real parts are flat at a specific non-Hermitian strength, which is denoted as “exceptional Dirac magic”. This flat band is characterized with a robust Chern number in moiré Brillouin zone. This is different from single-layer graphene in which Dirac/Weyl exceptional rings carrying opposite topological charges are coupled and annihilated.
@@
Note added-. Upon completion of our manuscript, we became aware of a recent paper on non-Hermitian twisted bilayer graphene, in which non-Hermiticity is on the hopping amplitude and topological aspects are undiscussed [61].
@@ Acknowledgements.
- CastroNeto et al. [2009] A. H. CastroNeto, F. Guinea, N. M. R. Peres, K. S. Novoselov, and A. K. Geim, The electronic properties of graphene, Reviews of modern physics 81, 109 (2009).
@@
- McCann [2006] E. McCann, Asymmetry gap in the electronic band structure of bilayer graphene, Phys. Rev. B 74, 161403(R) (2006).
@@
- Bistritzer and MacDonald [2011b] R. Bistritzer and A. H. MacDonald, Moiré butterflies in twisted bilayer graphene, Physical Review B 84, 035440 (2011b).
@@
- Luican et al. [2011] A. Luican, G. Li, A. Reina, J. Kong, R. R. Nair, K. S. Novoselov, A. K. Geim, and E. Andrei, Single-layer behavior and its breakdown in twisted graphene layers, Physical review letters 106, 126802 (2011).
@@
- Shao et al. [2024] K. Shao, H. Geng, E. Liu, J. L. Lado, W. Chen, and D. Y. Xing, Non-hermitian moiré valley filter, Physical Review Letters 132, 156301 (2024).
@@
- Król et al. [2022] M. Król, I. Septembre, P. Oliwa, M. Kkedziora, K. Łempicka-Mirek, M. Muszyński, R. Mazur, P. Morawiak, W. Piecek, P. Kula, et al., Annihilation of exceptional points from different dirac valleys in a 2d photonic system, Nature Communications 13, 5340 (2022).
@@
- Tarnopolsky et al. [2019a] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Phys. Rev. Lett. 122, 106405 (2019a).
@@
- Tarnopolsky et al. [2019b] G. Tarnopolsky, A. J. Kruchkov, and A. Vishwanath, Origin of magic angles in twisted bilayer graphene, Physical review letters 122, 106405 (2019b).
@@
- Esparza and Juricic [2024] J. P. Esparza and V. Juricic, Exceptional magic angles in non-hermitian twisted bilayer graphene, arXiv preprint arXiv:2408.08804 (2024).